# Flipping Out: Role of Arginine in Hydrophobic Polymer Collapse

### v2

@@ Abstract
We discovered that, with increasing concentration, the mechanism underlying the effects of arginine transitions from direct to indirect dominance. At 0.25 M, cavity formation and polymer-water interactions oppose polymer collapse, while arginine-polymer interactions favor collapse. The balance of these components gives rise to $\Delta\Delta G_{dir}>\Delta\Delta G_{ind}$, resulting in net stabilization of folded conformations and supporting the direct mechanism hypothesis (Fig. 2a).
In contrast, for the high-concentration regime (0.5 M and 1.0 M), cavity formation and polymer-water attractive interactions favor polymer collapse, while attractive arginine-polymer interactions favor extension of the hydrophobic polymer. In this case, indirect components dominated the free energy difference ($\Delta\Delta G_{dir}<\Delta\Delta G_{ind}$), stabilizing the polymer collapse and supporting the indirect hypothesis (Fig. 2a).
@@
While glycine is known to be an effective stabilizer of proteins,81, 82, 18 our observations obtained for guanidinium are somewhat surprising due to its common role as a protein denaturant 83, 84, 85. Several studies have stressed the importance of direct interactions in guanidinium-induced denaturation, primarily via breaking salt bridges, competing for intra-protein hydrogen bonds, and interacting with aromatic moieties via cation-pi stacking.84, 86, 87. Usually, this occurs at high concentrations of guanidinium salts. Our findings suggest that while guanidinium may stabilize hydrophobic interactions at low concentrations, this is outweighed by the denaturation mechanisms at high concentrations.
@@
Thermodynamic analyses of arginine effects on polymer collapse discussed above indicate that direct and indirect mechanisms co-exist and compete at all concentrations under study. To probe this further, we characterize the molecular interactions between arginine, water, and the polymer. Specifically, we look at hydrogen bonding between arginine and water to characterize arginine-water interactions. Preferential interactions are used to elucidate polymer-arginine-water interactions.
Fig. 3 describes hydrogen bonding interactions between arginine and water. Hydrogen bonds were calculated according to geometric criteria of a donor-acceptor distance of $\mathrm{r<=}$ 0.35 $\mathrm{nm}$ and an angle deviating less than 30o from 180o.88 Hydrogen bond existence correlation functions for water-water, guanidinium-water, carboxylate-water, and amine-water were estimated according to89, 90, 91
where $h_{ij}\left(t_{0}\right)$ is equal to 1 if there is a hydrogen bond between groups $i$ and $j$ at time $t_{0}$, and 0 if no hydrogen bond is present. An average over all possible values of the time origin $\mathrm{t_{0}}$ was taken over the last 5 ns of 20 ns equilibrium simulations.
Figure 3: Arginine-water interactions. (a) Hydrogen bonds observed between backbone groups (COO-, NH+${}_{3}^{+}$) and the guanidinium (Gdm+) sidechain with water. (b) Hydrogen bond existence correlation functions for water-water, COO--water, Gdm+-water, and NH+${}_{3}^{+}$-water. (c) Radial distribution function between OW and either CA (purple) or CZ (gold). (c, inset) Arrows denote trends observed with increasing arginine concentration. (d) A representative snapshot of hydrogen-bonding interactions involving arginine and water. Water molecules interacting with the Gdm+ sidechain are highlighted in yellow, while those interacting with NH+${}_{3}^{+}$ and COO- are shaded in blue and purple, respectively.
@@
Overall, the number of backbone-water hydrogen bonds was observed to be greater than sidechain-water hydrogen bonds (Fig. 3a). We further observed the fraction of occupied hydrogen bonding sites to be higher for backbone groups than the sidechain (Fig. S7). The hydrogen bond existence autocorrelation function for the 0.25 M arginine solution revealed hydrogen bonds formed between backbone-water atoms are, on average, longer-lived than sidechain-water hydrogen bonds (Fig. 3b). As arginine concentration was increased, hydrogen bond lifetimes were also observed to increase for both arginine-water and water-water interactions (Fig. S8).
@@
Pairwise radial distribution functions (RDFs) were computed between the water oxygen (OW) and either the alpha carbon (CA) or guanidinium carbon (CZ) of arginine to quantify the local structure of water around arginine molecules (Fig. 3c). The first peak in the OW-CA RDF was observed to increase slightly with concentration. This indicates preferential hydration of the backbone as more arginine molecules are introduced to the solution. There is, however, no such change observed in the OW-CZ RDF with concentration. A representative snapshot of the hydration shell of a single arginine molecule is shown in Fig. 3d (2D representation is shown in Fig. S9. Together, these results indicate that the backbone of arginine interacts favorably with water.
@@
While we found arginine preferentially interacts with water via its backbone, we hypothesized arginine interacts with the polymer via its sidechain. It has been reported elsewhere that dehydration of the guanidinium face is important in forming face-face stacking interactions in aqueous guanidinium solutions.92 In our case, the dehydrated face of guanidinium is expected to play a key role in direct arginine-polymer interactions, similar to interactions observed between guanidinium and hydrophobic/aromatic protein residues.93, 7, 20, 19
Distribution of arginine with respect to the polymer is described via the preferential interaction coefficient ($\Gamma_{PA}$),94, 95, 96
where $\mu$ is the chemical potential, $m$ is the concentration and $W$, $P$, and $A$ refer to water, polymer, and the additive, respectively. This parameter is calculated in simulations using the two-domain formula97, 98, 99 given by,
@@
Wyman-Tanford theory relates the dependence of any equilibrium process (such as protein folding) and preferential interaction as:100, 101, 102
$-\left(\frac{\partial\Delta G^{u}}{\partial\mu_{A}}\right)=\Gamma_{PA}^{u}-\Gamma_{PA}^{f}$ (6)
@@
As a result, denaturants are expected to have a greater preferential interaction coefficient in the unfolded ensemble, while stabilizing osmolytes have a greater preferential interaction coefficient in the folded ensemble41, 58, 103, 104. In 0.25 M arginine concentration, we observed greater preferential interactions with the folded state relative to the unfolded state (Fig. 4a; Fig. S11a-c). With increasing concentration, we observe a diminishing difference between $\Gamma_{PA}^{u}$ and $\Gamma_{PA}^{f}$, in line with the mechanistic flip from a direct- to indirect-dominated stabilization mechanism.
@@
Figure 4: Preferential interaction coefficient values as a function of the cut-off distance for the local domain of the hydrophobic polymer for (a) arginine, (b) guanidinium, and (c) glycine. Dashed lines indicate values for the unfolded state, while solid lines denote the folded state. Increasing concentration is denoted by increased shading (light to dark). Mean values and errors were estimated from three replicate simulations. Errors are reported as standard deviations from mean values.
@@
Preferential interaction coefficients for guanidinium and glycine solutions are shown in Fig. 4b and c, respectively. We observed that at all concentrations, guanidinium preferentially interacts with the hydrophobic polymer (Fig. 4b). This is consistent with experimental evidence, as well as a prior simulation study that observed attractive guanidinium-polymer interactions for a model hydrophobic polymer39, 70. Glycine, meanwhile, was found to be preferentially excluded from the local domain of the hydrophobic polymer (Fig. 4c). This finding is consistent with the observed preference for the backbone of arginine to hydrogen bond with water, relative to the sidechain. Elsewhere, glycine has been observed to deplete from the surface of several model miniproteins, consistent with our findings 104.
@@
To explore whether the preferential interactions of arginine with polymer and water are accompanied by preferential orientations, we computed an orientation parameter inspired by Shukla and Trout.105 This parameter was computed as the angle formed in three-dimensional space between polymer-CZ and CZ-CA vectors (Fig. 5a). The monomer closest to CZ is taken for the polymer-CZ vector. Angles where $\theta>90\degree$ indicate the arginine backbone orients towards the bulk solvent, while $\theta<90\degree$ indicates the arginine backbone orients towards the polymer. We observed that for all concentrations, the probability $P(\theta)$ is skewed towards angles greater than 90°$\degree$ (Fig. 5b-d).
@@
Further, for all arginine concentrations, the mean value of $\theta$ for the folded state is greater than that observed for unfolded configurations (Fig. 5b-d). This preferential orientation of arginine enables the hydrophobic face of the guanidinium sidechain to interact with the hydrophobic polymer, while extension of the backbone towards the bulk enables additional interactions with either water or other free arginine molecules. The greater ability for arginine to adopt preferred orientations in the folded state, particularly at 0.25 M, may partially explain the favorable $\Gamma_{PA}$ values described previously.
@@
Figure 5: Preferential orientation of arginine relative to the hydrophobic polymer. (a) Representation of the three-body angle, $\theta_{Pol-CZ-CA}$. $P(\theta)$ is shown for (b) 0.25 M, (c) 0.5 M, and (d) 1.0 M arginine concentrations. Solid lines denote the probability distribution for folded conformations, while dashed lines indicate the unfolded state. Vertical lines represent the ensemble average of each state.
@@
The simultaneous presence of competing direct and indirect effects implies that changes in the chemistry of a protein surface, the addition of co-additives to solution, or differences in sample preparation may cause significant changes in the mechanism of action of arginine. A shift towards the direct mechanism risks denaturing native proteins by disrupting electrostatic and hydrogen-bonding interactions. Conversely, a shift towards the indirect mechanism may yield stabilization of native proteins through preferential hydration. Our results suggest that arginine is uniquely situated for use in formulations due to its potential tunable, context-dependent properties. Hence, while arginine may not be considered the universal stabilizer it once was, its balance between direct- and indirect-driven stabilization of hydrophobic interactions solidifies its significance in formulation design.
{acknowledgement}

### v3

@@ 3.3 Mechanistic flip with increasing arginine concentration
We discovered that, with increasing concentration, the mechanism underlying the effects of arginine transitions from direct to indirect dominance. At 0.25 M, cavity formation and polymer-water interactions oppose polymer collapse, while arginine-polymer interactions favor collapse. The balance of these components gives rise to $\Delta\Delta G_{dir}>\Delta\Delta G_{ind}$, resulting in the net stabilization of folded conformations and supporting the direct mechanism hypothesis (Fig. 2a).
In contrast, for the high-concentration regime (0.5 M and 1.0 M), cavity formation and polymer-water attractive interactions favor polymer collapse, while attractive arginine-polymer interactions favor extension of the hydrophobic polymer. In this case, indirect components dominate the free energy difference ($\Delta\Delta G_{dir}<\Delta\Delta G_{ind}$), stabilizing polymer collapse and supporting the indirect hypothesis (Fig. 2a).
@@ 3.4 Distinct roles of arginine’s sidechain and backbone
While glycine is known to be an effective stabilizer of proteins,107, 108, 21 our observations obtained for guanidinium are somewhat surprising due to its common role as a protein denaturant.109, 110, 111 Several studies have stressed the importance of direct interactions in guanidinium-induced denaturation, primarily via breaking salt bridges, competing for intra-protein hydrogen bonds, and interacting with aromatic moieties via cation-pi stacking.110, 112, 113 Usually, this occurs at high concentrations of guanidinium salts. Our findings suggest that while guanidinium may stabilize hydrophobic interactions at low concentrations, this is outweighed by denaturing mechanisms at high concentrations.
3.5 Resolving polymer-arginine-water interactions
@@
Thermodynamic analyses of arginine effects on polymer collapse discussed above indicate that direct and indirect mechanisms co-exist and compete at all concentrations under study. To probe this further, we characterize the molecular interactions between arginine, water, and the polymer. Specifically, we look at hydrogen bonding between arginine and water to characterize arginine-water interactions. Preferential interactions are used to elucidate the balance of polymer-arginine-water interactions.
@@
To describe how arginine interacts with water, we considered hydrogen bonding interactions between water and backbone (COO-, NH+${}_{3}^{+}$) or sidechain (Gdm+) atoms of arginine. Overall, the number of backbone-water hydrogen bonds was observed to be greater than sidechain-water hydrogen bonds (Fig. 3a). We further observed the fraction of occupied hydrogen bonding sites to be higher for backbone groups than the sidechain (Fig. S7). The hydrogen bond existence autocorrelation function for the 0.25 M arginine solution revealed hydrogen bonds formed between backbone-water atoms are, on average, longer-lived than sidechain-water hydrogen bonds (Fig. 3b). As arginine concentration was increased, hydrogen bond lifetimes were also observed to increase for both arginine-water and water-water interactions (Fig. S8). In principle, this may be considered a stabilizing property of arginine, as a growing body of literature has supported the role of stabilizing osmolytes in increasing hydrogen bond lifetimes and reducing water dynamics.114, 115, 116, 117, 118, 119, 120
Figure 3: Arginine-water interactions. (a) Hydrogen bonds observed between backbone groups (COO-, NH+${}_{3}^{+}$) and the guanidinium (Gdm+) sidechain with water. (b) Hydrogen bond existence correlation functions for water-water, COO--water, Gdm+-water, and NH+${}_{3}^{+}$-water in 0.25 M arginine solution. (c) Radial distribution function between OW and either CA (purple) or CZ (gold). (c, inset) Arrows denote trends observed with increasing arginine concentration. (d) A representative snapshot of hydrogen-bonding interactions involving arginine and water. Water molecules interacting with the Gdm+ sidechain are highlighted in yellow, while those interacting with NH+${}_{3}^{+}$ and COO- are shaded in blue and purple, respectively.
@@
Pairwise radial distribution functions (RDFs) were computed between the water oxygen (OW) and either the alpha carbon (CA) or guanidinium carbon (CZ) of arginine to quantify the local structure of water around arginine molecules (Fig. 3c). The first peak in the OW-CA RDF was observed to increase slightly with concentration. This indicates preferential hydration of the backbone group as more arginine molecules are introduced to the solution. There is, however, no such change observed in the OW-CZ RDF with concentration. A representative snapshot of water interactions with a single arginine molecule is shown in Fig. 3d (2D representation is shown in Fig. S9). Together, these results indicate that the backbone of arginine is the primary site for interaction with water.
@@
While we found arginine preferentially interacts with water via its backbone, we hypothesized arginine interacts with the polymer via its sidechain. It has been reported elsewhere that dehydration of the planar guanidinium face is important in forming face-face stacking interactions in aqueous guanidinium solutions.121 In our case, the dehydrated face of guanidinium is expected to play a key role in direct arginine-polymer interactions, similar to interactions observed between guanidinium and hydrophobic/aromatic protein residues.122, 7, 23, 22
@@
Wyman-Tanford theory relates the dependence of any equilibrium process (such as protein folding) and preferential interaction as:123, 124, 125
$-\left(\frac{\partial\Delta G_{u}}{\partial\mu_{A}}\right)=\Gamma_{PA}^{u}-\Gamma_{PA}^{f}$ (7)
@@
where $\Gamma_{PA}^{u}$ represents the preferential interaction coefficient in the unfolded state, while $\Gamma_{PA}^{f}$ represents the folded state. As a result, denaturants are expected to have a greater preferential interaction coefficient in the unfolded ensemble, while stabilizing osmolytes have a greater preferential interaction coefficient in the folded ensemble.45, 62, 126, 127 In 0.25 M arginine concentration, we observed greater preferential interactions with the folded state relative to the unfolded state (Fig. 4a; Fig. S11a-c). With increasing concentration, we observe a diminishing difference between $\Gamma_{PA}^{u}$ and $\Gamma_{PA}^{f}$, in line with the mechanistic flip from a direct- to indirect-dominated stabilization mechanism.
@@
Figure 4: Preferential interaction coefficient values as a function of the cut-off distance for the local domain of the hydrophobic polymer for (a) arginine, (b) guanidinium, and (c) glycine. Dashed lines indicate values for the unfolded state, while solid lines indicate the folded state. Increasing concentration is denoted by increased shading (light to dark). Mean values and errors were estimated from three replicate simulations. Errors are reported as standard deviations from mean values.
@@
Preferential interaction coefficients for guanidinium and glycine solutions are shown in Fig. 4b and c, respectively. We observed that at all concentrations, guanidinium preferentially interacts with the hydrophobic polymer (Fig. 4b). This is consistent with experimental evidence, as well as a prior simulation study that observed attractive guanidinium-polymer interactions with a model hydrophobic polymer.43, 79 Glycine, meanwhile, was found to be preferentially excluded from the local domain of the hydrophobic polymer (Fig. 4c). This finding is consistent with the observed preference for the backbone of arginine to hydrogen bond with water, relative to the sidechain. Elsewhere, glycine has been observed to deplete from the surface of several model miniproteins, consistent with our findings. 127
@@
To explore whether the preferential interactions of arginine with polymer and water are accompanied by preferential orientations, we computed an orientation parameter inspired by Shukla and Trout.128 This parameter was computed as the angle formed in three-dimensional space between polymer-CZ and CZ-CA vectors (Fig. 5a). The monomer closest to CZ is taken for the polymer-CZ vector. Angles where $\theta>90\degree$ indicate the arginine backbone orients towards the bulk solvent, while $\theta<90\degree$ indicates the arginine backbone orients towards the polymer. We observed that for all concentrations, the probability $P(\theta)$ is skewed towards angles greater than 90°$\degree$ (Fig. 5b).
@@
Further, for all arginine concentrations, the mean value of $\theta$ for the folded state is greater than that observed for unfolded configurations (Fig. 5b). This preferential orientation of arginine enables the hydrophobic face of the guanidinium sidechain to interact with the hydrophobic polymer, while extension of the backbone towards the bulk enables additional interactions with either water or other free arginine molecules. The greater ability for arginine to adopt preferred orientations in the folded state, particularly at 0.25 M, may partially explain the favorable $\Gamma_{PA}$ values described previously.
@@
Figure 5: Preferential orientation of arginine relative to the hydrophobic polymer. (a) Representation of the three-body angle, $\theta$. (b) $P(\theta)$ is shown for 0.25 M, 0.5 M, and 1.0 M arginine concentrations. Solid lines denote the probability distribution for folded conformations, while dashed lines indicate the unfolded state. Increased concentration is denoted by increased shading.
4 Contextualizing the Mechanistic Flip of Arginine
4.1 Arginine destabilizes a charged polymer
To explore whether we could shift arginine further towards a direct-dominated mechanism, we modified our hydrophobic polymer to include four beads with opposing partial charges (Fig.  6a). In pure water, charged polymer collapse was observed to be favorable, resulting in a prominent free energy minimum at $R_{g}\sim$ 0.5 nm (Fig.  6b). Relative to the hydrophobic polymer, this is indicative of a conformational preference of the polymer to adopt hairpin-like, rather than globule-like, configurations (I and II in Fig.  6a, respectively).
Upon addition of either 0.25 M or 1.0 M arginine, folding of the charged polymer becomes less favorable (Fig.  6c). We attribute this to a preference for arginine to interact with the charged sites, which are more accessible to arginine in the unfolded state. Indeed, at both concentrations, attractive polymer-arginine interactions dominate (Fig.  6d), driving polymer unfolding. This model demonstrates that even subtle changes to the chemistry of a macromolecule can re-balance arginine mechanisms.
Figure 6: Charged polymer simulations. (a) Locations of the negative (blue) and positive (red) partial charges. (b) PMFs associated with the charged polymers. (c) Free energy of polymer collapse as a function of arginine concentration. (d) Change in PMF components in Arg solution relative to pure water, for the charged polymer. Increasing concentration is denoted by increased shading.
4.2 Implications for formulation design
Arginine-induced destabilization of a charged polymer illustrates that, when hydrophobic interactions compete with other effects, the effectiveness of arginine as a stabilizer can be altered. To better understand this feature of arginine, we explored the temperature stability of two models for formulation design: PPV and HEWL.
PPV is a non-enveloped virus with a single stranded DNA genome.129 The PPV capsid is a spherical shell comprised of 60 copies of viral proteins (VP) VP1, VP2, and VP3 in a 1:10:1 ratio, arranged in an icosahedral symmetry.92 Due to the relatively small size and structural simplicity of PPV, recent studies have employed the virus as a model for investigating virus purification and thermostabilization techniques.94, 130, 131, 95
HEWL, meanwhile, is a relatively small protein with a well-defined fold (PDB: 1LYZ) and high stability.93, 132 Due to these features, it has been widely used as a model for protein folding133, 134, 135 and exploring osmolyte effects.136, 137, 107 HEWL has also been used to investigate aggregation-suppressing effects of arginine. Several studies have proposed that arginine interacts favorably with aromatic and acidic residues of HEWL, which limits solvent exposure of aggregation-prone patches.7, 22, 23, 20, 138 Arginine has also been observed to enhance the heat-induced aggregation of bovine serum albumin and $\beta$-lactoglobulin, but not HEWL – highlighting its context-dependent effects.11
To understand the effect of arginine on these biomolecules, temperature stability assays were carried out at different concentrations of arginine. For PPV, we completed an infectivity assay (see SI for experimental details) following virus incubation at 60 °C for 3 days – sufficiently long to observe a significant decrease in the infectious titer.94 Log reduction values (LRV) describe the decrease in the infectious titer of heat-treated virus, relative to the initial virus solution. We found, at all investigated arginine concentrations, more infectious PPV is lost relative to in buffer alone, resulting in negative $\Delta$LRV values (Fig.  7a). Such a finding indicates reduced temperature stability of the virus in arginine solutions.
Figure 7: Temperature stability of (a) PPV and (b) HEWL as a function of arginine concentration. $\Delta$LRV is reported as LRVBuffer - LRVArg, while $\Delta$THE $=$ THE,Arg - THE,buffer. Increasing concentration is denoted by increased shading. Mean and standard deviation are estimated from triplicate measurements.
For HEWL, we completed a thermal shift assay (see SI for details) to quantify the hydrophobic exposure temperature ($T_{HE}$) of the protein.96, 97 At low concentrations (0.2-1.0 M) of arginine, a decrease in $T_{HE}$ is observed relative to buffer alone, indicating destabilization of HEWL (Fig.  7b). With increasing arginine concentration (1.5 M), the temperature stability of HEWL improves. Such a finding closely resembles the temperature stability of ovalbumin and lysozyme in arginine solutions reported by Vagenende et al.20 Additionally, 1.0 M arginine was found to increase protein interactions with hydrophobic chromatographic materials.20 This finding highlights that, when experimentally-observed effects of arginine mirror what we observed with the hydrophobic polymer, hydrophobic interactions are likely to dominate. On the other hand, our results contrast findings presented by Platts and Falconer,21 who observed temperature-stabilizing effects at low arginine concentrations and destabilization of HEWL at high concentrations.
4.3 Molecular-level investigation of formulation models
The lack of consensus regarding the effects of arginine – further accentuated by our investigation of PPV and HEWL temperature stability – may be rationalized by the positioning of arginine at the edge of the mechanistic flip described in the present work. Additionally, we have shown that when charged beads are added to a hydrophobic polymer, this subtle change results in complete arginine-induced destabilization. To better understand the connection between the mechanistic flip of arginine and the relevance of differing molecular contexts, we completed straightforward MD simulations of HEWL and PPV systems in arginine solution.
The PPV major capsid protein, VP2, is known to self-assemble into virus-like particles that are non-pathogenic, but are morphologically similar to native PPV virions.139 Experimental identification of this structure92 enables PPV for molecular-level investigations via molecular simulations. On the capsid surface, prominent structural features include a canyon surrounding a pore at the 5-fold axis of symmetry, protruding spikes located at the 3-fold axis of symmetry, and a dimple on the 2-fold axis of symmetry.92 We found that the use of 15 VP2 proteins is the minimum system size required to capture these structural features, enabling simulations at timescales accessible to MD without sacrificing atomistic resolution. Despite this truncation of the viral capsid, our resulting system size was $\sim$750,000 atoms.
At the PPV surface in 0.25 M arginine solution, we observed preferential arginine accumulation at various sites across the surface (Fig.  8a,c). Specifically, via calculation of contact coefficients of different residue types distributed across the PPV capsid surface, we identified significant accumulation of arginine near negatively charged glutamate and aspartate residues. From these findings, we hypothesize that direct interactions between arginine and charged residues at the PPV surface drive the instability observed from experiments. In this case, the dominating mechanism may be similar to that observed for arginine-induced unfolding of a charged polymer.
For HEWL, we completed simulations of 0.1 M, 0.2 M, 1.0 M, and 2.0 M arginine solutions. With increasing concentration, we observed an increase in preferential exclusion of arginine and reduction of arginine density local to the protein (Fig.  8b,d). This finding, along with the increase in HEWL melting temperature observed at high arginine concentration, suggests that arginine imparts stabilization of HEWL via indirect effects. On the other hand, direct arginine-HEWL interactions may be destabilizing, which explains the decreased melting temperature of HEWL while arginine is less excluded from the local domain of the protein.
Figure 8: Arginine distributions from PPV and HEWL systems. (a) Arginine contact coefficients at the PPV 5-fold surface. Bars are shown for different residue types, hydrophobic (green), hydrophilic (yellow), positive (blue), and negative (red), and was computed separately for backbone atoms (gray). (b) Preferential interaction coefficient for arginine/HEWL. Increasing concentration is denoted by increased shading (light to dark). (c) Representative snapshot and volumetric density plots of the (c) PPV 5-fold surface and (d) HEWL in arginine solutions. In c and d, PPV and HEWL are shown in a surface representation. Protein residues are colored according to residue type as in panel a. Arginine density at an isovalue of 0.1 atoms / Å3 is shown in black mesh representation.
5 Conclusions
@@
The simultaneous presence of competing direct and indirect effects implies that changes in the chemistry of a protein surface, the addition of co-additives to solution, or differences in sample preparation may cause significant changes in the mechanism of action of arginine. A shift towards the direct mechanism risks guanidinium-like denaturation of native proteins by disrupting electrostatic and hydrogen-bonding interactions. Conversely, a shift towards the indirect mechanism may yield glycine-like stabilization of native proteins through preferential hydration. Indeed, we illustrated examples of the context-dependent effects of arginine through models of protein (HEWL) and virus (PPV) stability. We observed that arginine destabilizes PPV across a wide concentration range, which we attribute to destabilizing direct arginine-PPV interactions. This mechanism resembles the effects of arginine on a charged polymer model, suggesting that in both cases, arginine-induced destabilization via charge-charge interactions outweighs arginine-induced stabilization of hydrophobic interactions. We further found that arginine destabilizes HEWL at low concentrations and stabilizes HEWL at high concentrations, which we found is associated with an increased exclusion of arginine from the HEWL surface. This mechanism closely resembles the concentration-dependence of the indirect mechanism of arginine in a hydrophobic polymer model. Hence, at high concentrations, HEWL destabilization due to arginine charge-charge interactions may be opposed by an overall stabilization of hydrophobic interactions.
Together, our results suggest that arginine is uniquely situated for use in formulations due to its tunable, context-dependent properties. Hence, while arginine may not be considered the universal stabilizer it once was, its balance between direct- and indirect-driven stabilization of hydrophobic interactions solidifies its significance in formulation design.
Author Contributions
J.Z.: Investigation - Computational, Validation, Formal Analysis, Visualization, Writing - Original Draft, Writing - Review and Editing; P.M.: Investigation - Computational, Validation, Writing - Review and Editing; I.T.: Conceptualization, Investigation - Experimental, Writing - Review and Editing; X.Z.: Investigation - Experimental, Writing - Review and Editing; C.H.: Conceptualization, Funding Acquisition, Resources, Writing - Review and Editing; S.P.: Conceptualization, Funding Acquisition, Resources, Writing - Review and Editing; S.S.: Conceptualization, Methodology, Validation, Supervision, Project Administration, Funding Acquisition, Resources, Writing - Review and Editing
Conflicts of interest
There are no conflicts to declare.
Acknowledgements